OpenStudy (anonymous):

Sketch the region bounded by the curves and find its area. x+y-y^3=0 , x-y+y^2=0 how can you tell which to use? in terms of x or y? (please explain in village idiot terminology b/c my algebra is weak!!! I am trying to improve it lol. calculus is easy, algebra is not :/ )

OpenStudy (anonymous):

either or. you can solve this for x or y

OpenStudy (anonymous):

since x is by itself i think that'd be the easiest however you'd have to switch your graphing calculator to graph the curves

OpenStudy (anonymous):

without a calculator :o and how do you get those to be in terms of x ? y=.....Myt algebra deficient brain is just not seeing it.

OpenStudy (anonymous):

to be honest i really think you shouldn't since it much easier to find when x=

OpenStudy (anonymous):

\[x=y^3-y\] \[x=y-y^2 \]

OpenStudy (anonymous):

when y = 0 both = 0 so that can be one of your your limits

OpenStudy (anonymous):

haha wait thats all you do ! -_-

OpenStudy (anonymous):

yes then you integrate in respects to y (dy

OpenStudy (anonymous):

now you have to find when these two intersect

OpenStudy (anonymous):

\[y^3-y=y-y^2 \] \[y^3 +y^2 -2y=0\] \[y(y^2+y-2)=0\]

OpenStudy (anonymous):

there is three intersectiosn so you're going to need two limits =]

OpenStudy (anonymous):

i mean two integrations

OpenStudy (anonymous):

yeah. adding them. so i use quadratic equation two find other two intersections right

OpenStudy (anonymous):

they factor (x+2)(x-1)

OpenStudy (anonymous):

(y+2)(y-1)*

OpenStudy (anonymous):

oh derp. -.- i see now. that was simpler than i imagine it. i got it from here. thank you !

OpenStudy (anonymous):

also since the area of one of the regions is below the x axis... it's going to come uot negative... which will create a net, you might want to htink about that